OpenStudy (anonymous):

How many three different digit numbers can be formed using the digits 2,3,5,6,7. ASAP with process

OpenStudy (zzr0ck3r):

can we repeat?

OpenStudy (anonymous):

We have to find how many of 3-digit numbers can be formed by using 2,3,5,6,7. When every digit should be different in the number

OpenStudy (anonymous):

In the first digit you can have 5 different numbers (all the digits), then in the second you can have 5-1=4 possible numbers because it cannot be equal to the first digit. Then in the third digit you are left with only 4-1=3 possible digits because two of them have been already used in the first two places. Therefore 5x4x3=60 possible combinations.

OpenStudy (madhu.mukherjee.946):

Using permutation you can solve this problem as - First place can be filled with 5 ways Second place can be filled with 4 ways Third place can be filled with 3 ways So ,answer is 5*4*3=60